# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v2

@@ V-A Proof of Lemma 1
where $\bm{\delta}\left(\mathbf{w}\right)$ is a $d$-dimensional function of $\mathbf{w}$ denoting the model distortion caused by bit errors. Under Assumption 4, each entry of $\bm{\delta}\left(\mathbf{w}\right)$ depends only on $w_{d}$, i.e., the $d$-th dimension of $\mathbf{w}$. Denote $\delta\left(w_{d}\right)$ as the $d$-th dimension of $\bm{\delta}\left(\mathbf{w}\right)$. Therefore, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
@@
The expected bias of $w_{d}$ is bounded as follows:
$\begin{split}|\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]-w_{d}|&\leq b\left(1-b\right)^{N-1}\left(2^{N}-1\right)\Delta\\ &=b\left(1-b\right)^{N-1}M.\end{split}$ (21)
The mean square error of $w_{d}$ is given as follows:
$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M^{2}.\end{split}$ (22)
@@
$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]&=\sum_{j=1}^{d}\mathbb{E}\left[||\tilde{w}_{j}-w_{j}||^{2}|\mathbf{w}\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M^{2}.\end{split}$ (23)
@@ V-B Proof of Lemma 2
We define
$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}),$ (24)
$\mathbf{\bar{w}}_{r,e}$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k},$ (25)
for each communication round $r=0,...,R-1$ and each local iteration $e=0,\cdots,E-1$, where $\mathbf{w}_{r+1}$ denotes the updated global model and $\mathbf{\bar{w}}_{r,e}$ is the average of client models after the $e$-th local iteration.
@@
For $\mathbf{\bar{w}}_{r,E}$ and $\mathbf{w}_{r+1}$ defined in (24) and (25), we have
@@
The proof can be obtained by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ to the LHS of (26). ∎
@@
The proof can be obtained using the $L$-smoothness of $f(\cdot)$, which is omitted for brevity. ∎
@@
$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (28)
@@
Lemma 28 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}$. By applying Lemmas 26 and 28 to Lemma 27, we obtain the following inequality:
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (29)
By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ holds for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (29) is derived as follows:
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{n}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (30)
where (a) is due to the $L$-smoothness of $f(\cdot)$ and (b) is due to the local updating rules. Based on the Assumptions 2 and 3, we have
@@
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{\leq}\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2n}\\ &\ \ \ +\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2},\end{split}$ (32)
@@
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2n}\left[1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (33)
@@
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}\\ &+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (34)
Summing (34) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we obtain the following expression:
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{L(L+2)}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{1}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.\end{split}$ (35)
Based on (23) and (21) respectively, we have
$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}\right]&=\mathbb{E}\left[\mathbb{E}\left[||\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}||^{2}|\mathbf{w}_{r}\right]\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b_{r}\left(1-b_{r}\right)^{N-1}\cdot M_{r}^{2},\end{split}$ (36)
$\begin{split}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}&=\sum_{j=1}^{d}\mathbb{E}\left[|\delta(w_{r,d})|^{2}|{w}_{r,d}\right]\\ &\leq\sum_{j=1}^{d}\mathbb{E}\left[(b_{r}(1-b_{r})^{N-1}M)^{2}|{w}_{r,d}\right]\\ &\leq d(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}.\end{split}$ (37)
By substituting (36) and (37) into (35), and denote $f^{*}$ as the minimum global training loss, we have
@@
$+\frac{L(L+2)d(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
$+\frac{d}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}M_{r}^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (38)
Denoting $\hat{M}=max(M_{0},M_{1},\cdots,M_{R-1})$, we have
$+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}$
$+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}.$ (39)

### v3

@@ V-A Proof of Lemma 1
where $\bm{\delta}\left(\mathbf{w}\right)$ is a $D$-dimensional random function of $\mathbf{w}$ denoting the model distortion caused by bit errors. Under Assumption 4, each entry of $\bm{\delta}\left(\mathbf{w}\right)$ depends only on $w_{d}$, i.e., the $d$-th dimension of $\mathbf{w}$. Denote $\delta\left(w_{d}\right)$ as the $d$-th dimension of $\bm{\delta}\left(\mathbf{w}\right)$. Therefore, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
@@
where $\Delta=\frac{M_{\mathbf{w}}}{2^{N-1}}$. Note that the second term on the RHS of (21) is the non-zero expected bias caused by random bit errors. Thus, the mean square error of $w_{d}$ is given as follows:
$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M_{\mathbf{w}}^{2}.\end{split}$ (22)
@@
$\begin{split}\mathbb{E}\left[||\bm{\delta}(\mathbf{w})||^{2}|\mathbf{w}\right]&=\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]\\ &=\sum_{d=1}^{D}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|\mathbf{w}\right]\\ &=\frac{D(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M_{\mathbf{w}}^{2}.\end{split}$ (23)
@@ V-B Proof of Lemma 2
We rewrite (5) by substituting $\mathbf{w}_{r,0}^{k}=\tilde{\mathbf{w}}_{r}$ as
$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{K}\sum_{k=1}^{K}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}).$ (24)
@@
Based on the definitions of $\mathbf{w}_{r+1}$ and $\mathbf{\bar{w}}_{r,E}$, we have
@@
The proof is straightforward by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ in (24) to the LHS of (25). ∎
@@
The proof can be obtained using the $L$-smoothness property of $f(\cdot)$, which is omitted for brevity. ∎
@@
$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2K}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (27)
@@
Lemma 27 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}$.
By applying Lemmas 25 and 27 to Lemma 26, we obtain the following inequality:
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2K}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (28)
By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (28) is derived as follows:
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{K}\sum_{k=1}^{K}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (29)
where (a) is due to the $L$-smoothness property of $f(\cdot)$ and (b) applies the local updating rules in (3). Based on the Assumptions 2 and 3, we have
@@
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2K^{2}}\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\leq\frac{L^{2}\eta^{2}}{2K^{2}}\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2K}\\ &\ \ \ +\frac{(L^{2}+1)\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}.\end{split}$ (31)
@@
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{(L+1)^{2}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2K}\left[1-\frac{K+L}{K}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{K}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (32)
@@
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{(L+1)^{2}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2K}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(K+1)E(E-1)}{2K}.\end{split}$ (33)
Summing (33) over $r=0,\cdots,R-1$, rearranging the terms, and multiplying $\frac{2}{\eta RE}$ on both sides, we have
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-\mathbb{E}f(\mathbf{w}_{R}))}{\eta RE}\\ &+\frac{(L+1)^{2}}{\eta RE}\sum_{r=0}^{R-1}\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(K+1)(E-1)}{K}.\end{split}$ (34)
By substituting (23) into (34), and denote $f^{*}$ as the minimum global training loss, we have
@@
$+\frac{(L+1)^{2}D(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}M_{r}^{2}$
$+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(K+1)(E-1)}{K}.$ (35)
Finally, based on Assumption 6, we are able to show (9) in Lemma 2.